OpenStudy (erinweeks):

Simplify.∜400/∜5 A.5∜2 B. 2∜5 C.2√5 D. ∜80

OpenStudy (erinweeks):

@jim_thompson5910

OpenStudy (tkhunny):

\(\dfrac{\sqrt[4]{400}}{\sqrt[4]{5}} = \sqrt[4]{\dfrac{400}{5}} = \sqrt[4]{80} = \sqrt[4]{16\cdot5}\) Can you finish?

OpenStudy (erinweeks):

so it would be D correct?

OpenStudy (tkhunny):

Well, if you want to stop in the middle and entirely ignore my last step and the step after that - which I encouraged you to do.

OpenStudy (erinweeks):

well 16 * 5 = 80 so it would be d?

OpenStudy (tkhunny):

We had 80 and then we had 16*5. That is a hint as to the direction we should go. Please don't go back the other direction. Does this look like we are doing anything? 80 = 16*5 = 80 Hint: \(16 = 2^{4}\)

OpenStudy (erinweeks):

im confused lbvs

OpenStudy (tkhunny):

So far, you haven't shown any work at all. You MUST demonstrate that you have SOME connection to the work at hand. \(\sqrt[4]{16\cdot5} = \sqrt[4]{2^{4}\cdot5} = \sqrt[4]{2^{4}}\cdot\sqrt[4]{5}\) Now, YOU need to finish. One more thing to do.